OpenStudy (anonymous):

Find the y value for the point C such that AC and BC form a 2:3 ratio. Segment AB is shown. A is at negative 3, 5. B is at 3, 0. 1.2 3 2.4 2

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

2

OpenStudy (anonymous):



OpenStudy (anonymous):

That means \(\vec c =\dfrac{2}{5}\vec a+ \dfrac{3}{5}\vec b\) and \(\vec a= \left(\begin{matrix}-3\\5\end{matrix}\right)\) \(\vec b= \left(\begin{matrix}3\\0\end{matrix}\right)\) so, \(\vec c =\dfrac{2}{5}\left(\begin{matrix}-3\\5\end{matrix}\right)+ \dfrac{3}{5}\left(\begin{matrix}3\\0\end{matrix}\right)=\left(\begin{matrix}3/5\\2\end{matrix}\right)\) which show y =2